OpenStudy (anonymous):

F'(x) for F(x)=int(ln(1+t^2)) from -x^2 to x^2?

OpenStudy (john_es):

You have, \[F(x)=\int^{x^2}_{-x^2}\ln(1+t^2)dt=\int^{\alpha(x)}_{-\alpha(x)}f(t)dt\] So \[F'(x)=f(x^2)\alpha'(x)-f(-x^2)(-\alpha'(x))\] Try it. It is an application of the well known theorem, http://en.wikipedia.org/wiki/Fundamental_theorem_of_calculus

OpenStudy (anonymous):

thank you, originally i ended up with all terms canceling but i now i have \[8x ^{3}\ln(1+x^{4})\] i hope this is correct?

OpenStudy (john_es):

Nearly this. I would say, \[F'(x)=4x\ln(1+x^4)\] As \[\alpha'(x)=2x\]

OpenStudy (anonymous):

thank you so much!